OpenStudy (anonymous):

You decide to buy a 60 in plasma-screen TV. You borrow $4300 from the electronics store at 1.8% per month simple interest, with no payments due for 2 years from the date of purchase. How much will you owe on the loan after 15 months? Round your answer to the nearest whole dollar.

OpenStudy (anonymous):

f(x) = 4300 + 4300(0.18)(x) f(15) = 4300 + 4300(0.18)(15)

OpenStudy (radar):

Use caution in converting the percentage of 1.8% to its decimal equivalent.

OpenStudy (shamim):

i think the rate of simple interest will b 1.8%=1.8/100=0.018